OpenStudy (anonymous):

remove the radical from the denominator below. do not simplify the answer. -2(square root of 15)/6-(square root of 13)

OpenStudy (loser66):

\[\frac{ -2\sqrt{15} }{ 6-\sqrt{13} }\]

OpenStudy (loser66):

is it yours? any idea?

OpenStudy (anonymous):

\[-2\sqrt{15}\div(-6\sqrt{13})\]

OpenStudy (anonymous):

how do i do this

OpenStudy (dan815):